# Quantum Machine Learning and Grover’s Algorithm for Quantum Optimization of Robotic Manipulators

### v1

@@ Abstract
Optimizing high-degree-of-freedom robotic manipulators requires searching complex, high-dimensional configuration spaces, a task that is computationally challenging for classical methods. This paper introduces a quantum-native framework that integrates Quantum Machine Learning (QML) with Grover’s algorithm to solve kinematic optimization problems efficiently. A parameterized quantum circuit is trained to approximate the forward kinematics model, which then constructs an oracle to identify optimal configurations. Grover’s algorithm leverages this oracle to provide a quadratic reduction in search complexity. Demonstrated on 1-DoF, 2-DoF, and dual-arm manipulator tasks, the method achieves significant speedups—up to 93x over classical optimizers like Nelder-Mead—as problem dimensionality increases. This work establishes a foundational, quantum-native framework for robot kinematic optimization, effectively bridging quantum computing and robotics problems.
Index Terms:
Quantum Circuit Learning, Grover’s Algorithm, Robotic Manipulators, Kinematic Optimization, Quantum Machine Learning, Quantum Computing
@@ I Introduction
Robotic manipulators are pivotal in modern automation, manufacturing, and precision tasks across industries such as automotive assembly, medical surgery, and aerospace [1, 2, 3]. These systems require precise end-effector positioning, achieved through solving the inverse kinematics (IK) problem, which determines joint configurations to reach a target pose [4, 5]. Classical optimization methods for IK, including gradient-based techniques and exhaustive searches, perform adequately for manipulators with low degrees of freedom (DoF) and simple workspaces. However, they encounter limitations in high-dimensional, nonlinear manifolds [6, 7]. For instance, in closed-loop architectures or when comparing multiple design candidates, the challenge intensifies due to nonlinear constraints, high-dimensional spaces, and mixed discrete-continuous parameters [8]. Exhaustive searches over discretized joints scale exponentially as $O(k^{N})$ for an $N$-DoF system with $k$ values per joint, while advanced algorithms exhibit polynomial scaling, rendering them intractable for high-DoF manipulators. This computational bottleneck underscores the necessity for quantum computing, which leverages superposition and entanglement to enable parallel processing and asymptotic advantages in search and optimization [9, 10, 11]. Grover’s algorithm, for example, searches an unsorted database of $M$ elements in $O(\sqrt{M})$ queries, offering a quadratic speedup over classical $O(M)$ methods [12]. In high-DoF IK, where $M\sim k^{N}$, this translates to efficiency gains for exploring constrained configuration spaces. The advantages of quantum approaches—such as handling exponential complexity and providing global optima in non-convex landscapes—make them essential for advancing robotics beyond classical limits, particularly as quantum hardware matures.
@@
Recent quantum applications in robotics include Grover’s algorithm for optimal control in six-jointed arms [13], reformulating control as quantum search with simulations, and quantum annealing for IK as quadratic unconstrained binary optimization (QUBO) problems, achieving over 30-fold speedups in large instances [14]. Complementary efforts explore QML for robot test oracles [15], quantum-inspired reinforcement learning (RL) for control [16], and Grover-based schemes for classification tasks [17]. Additional works apply quantum-inspired sliding-mode control to enhance precision in articulated arms [18] and quantum deep RL for navigation [19]. These developments highlight quantum’s versatility but reveal a gap: no prior integration of QML with Grover’s for kinematic optimization in a quantum-native approach. To address this gap, we propose a quantum framework integrating QML—employing parameterized circuits to learn kinematic patterns—with Grover’s algorithm for efficient search. Leveraging hardware progress, such as Google’s Willow processor for scalable, error-corrected systems [20], this approach enables handling of higher-dimensional robotic problems. The framework comprises a QML block for state preparation and a Grover block for optimization (Fig. 1), where an oracle marks low-cost states based on observables (Table I). The QML circuit $U_{\text{QML}}(\psi)$ trains to approximate forward kinematics, generating a structured superposition as Grover’s input. The oracle, built from QML, applies phase flips to mark solutions, with diffusion amplifying probabilities iteratively. This leverages quantum parallelism for simultaneous configuration evaluation, accelerating optimal solution identification. Cost functions incorporate observables like position error, encoded as diagonal Hamiltonians for parallel quantum assessment. Table I lists representative examples.
@@
Note: Observables are diagonal in the computational basis, enabling quantum-parallel evaluation. $\bm{p}_{k}$, $\bm{p}_{\text{target}}$: end-effector and target positions; $\bm{R}_{k}$, $\bm{R}_{\text{target}}$: orientation matrices, $d_{R}$: SO(3) geodesic; $\bm{J}_{k}$: Jacobian; $\bm{\tau}_{k}$: torques; $d_{\text{min}}(\bm{q}_{k})$: minimum obstacle distance for joints $\bm{q}_{k}$. The projector $\lvert k\rangle\langle k\rvert$ maps configuration $k$ onto its quantum state.
@@
By recasting classical IK into quantum states and operations, this framework reduces evaluations for large spaces. Our primary contribution is formulating robotic optimization in a quantum-native manner, solved via Grover integrated with QML. The paper is organized as follows: Section II reviews related work; Section III details the framework; Section IV describes case studies; Section V presents evaluations; Section VI concludes.
II Background and Related Work
@@
This section surveys quantum computing advancements in robotics and allied fields, emphasizing motivations for our framework. For accessibility to robotics audiences, we briefly explain key quantum concepts: qubits, superposition, entanglement, and bra-ket notation. We define acronyms on first use where necessary.
II-A Robotic Manipulators: Fundamentals and Challenges
Robotic manipulators, typically serial chains of rigid links connected by revolute or prismatic joints, enable precise motion in 3D space. Common architectures include articulated arms (e.g., 6-DoF industrial robots like KUKA KR or ABB IRB) and parallel manipulators for enhanced stiffness and dynamics [3]. Kinematics governs their motion: forward kinematics (FK) maps joint variables $\bm{q}\in\mathbb{R}^{N}$ to end-effector pose $\bm{x}\in SE(3)$, often via Denavit-Hartenberg parameters or position closure [21]; IK inverts this, solving $\bm{q}=f^{-1}(\bm{x})$, which is nonlinear and multi-valued for redundant systems ($N>6$). The problem becomes complex when optimizing a robot’s geometric and performance parameters for task-specific suitability, as this involves FK, IK, velocity, and dynamic modeling. Classical solvers—geometric for low-DoF, iterative (e.g., Newton-Raphson) or metaheuristic (e.g., PSO, GA) for higher—face scalability issues: local minima, singularities (where Jacobian $\bm{J}(\bm{q})$ loses rank), and exponential complexity in high-DoF or constrained environments [5]. These challenges motivate quantum paradigms, which exploit parallelism for global optimization.
II-B Quantum Computing Essentials for Robotics
Quantum computers operate on qubits, whose states $\lvert\psi\rangle=\alpha\lvert 0\rangle+\beta\lvert 1\rangle$ ($\alpha,\beta\in\mathbb{C}$, with $|\alpha|^{2}+|\beta|^{2}=1$ giving measurement probabilities) enable superposition—representing multiple robotic configurations simultaneously. Entanglement correlates qubits, facilitating joint operations. In bra-ket notation, $\lvert\cdot\rangle$ denotes column vectors, $\langle\cdot\rvert$ their conjugate transposes; projectors like $\lvert k\rangle\langle k\rvert$ isolate states for observables. Quantum circuits apply unitary gates (e.g., Hadamard for superposition, CNOT for entanglement) on qubit registers. Algorithms like Grover’s amplify desired states via oracles and diffusion, ideal for search in robotic configuration spaces [12]. QAOA approximates combinatorial optima [22]. These offer quantum advantage for NP-hard problems in robotics. For foundational concepts, refer to Parts I and II of [23].
II-C Quantum Applications in Robotics and Related Domains
@@
Quantum methods have emerged to address robotics challenges. Dahassa et al. [13] reformulated optimal control as Grover’s search for a six-jointed arm, using quantum comparators and sigmoid updates for convergence, simulated classically. Salloum et al. [14] cast IK as QUBO for quantum annealing, achieving 30x speedups via hybrid solvers on D-Wave hardware. In QML, Wang et al. [15] used quantum reservoir computing for robot test oracles, reducing errors by 15%. Quantum-inspired RL aids control: Yu et al. [16] enhanced quantum system control in noisy environments; Kirchner et al. [19] applied quantum deep RL to navigation. Grover-based learning supports classification: Du et al. [17] improved schemes with fewer measurements; related to robotics via sensor data processing [24]. Other works include quantum teleportation for remote robot control [25] and hybrid LLM-Q-learning for adaptive robotics [26]. Table II summarizes the mapping between application domains and quantum algorithms, highlighting robotics’ emerging role.
@@
These developments motivate our QML-Grover integration, bridging learning and search for kinematic optimization. The following subsection overviews the proposed framework to provide a smooth transition to the methodology.
II-D Overview of the Proposed Framework
Our framework encodes kinematic parameters $\bm{z}\in\mathbb{R}^{d}$ into a parameterized circuit $U_{\text{QML}}(\bm{\theta})$ approximating forward kinematics $f(\bm{z})$. Data encoding uses rotation gates ($R_{X},R_{Y},R_{Z}$); training minimizes cost $L(\bm{z})=\|f(\bm{z})-\bm{p}_{\text{target}}\|^{2}$. For Grover’s, discretize $\bm{z}$ with $n_{i}$ qubits per parameter, yielding $N=\sum n_{i}$ qubits and space $\mathbb{C}^{2^{N}}$. Basis states $\lvert x\rangle$ map to configurations $\bm{z}_{x}$. The oracle marks low-cost states; Grover amplifies them, yielding near-optimal solutions. This integration exploits QML’s representation and Grover’s efficiency in a quantum-native approach.
@@ III-A Problem Statement: Robotic Kinematics Optimization
When geometric and joint parameters are unknown but the task is specified, optimization identifies parameters $\bm{d}$ and joint configuration $\bm{q}$ enabling task achievement. For each candidate $\bm{d}$, inverse kinematics checks existence of $\bm{q}$ achieving $T_{\text{target}}$. If infeasible or error exceeds tolerance, $\bm{d}$ is updated iteratively. This combines IK with parameter adjustment, ensuring the final design achieves $T_{\text{target}}$ within kinematic limits. Formally: Consider an $n$-DoF manipulator with $\bm{q}=[q_{1},\dots,q_{n}]^{\top}$, forward kinematics $f_{\text{FK}}:\mathbb{R}^{n}\times\mathcal{D}\rightarrow SE(3)$ parameterized by $\bm{d}\in\mathcal{D}$, and target pose $T_{\text{target}}=[\bm{R}_{\text{target}},\bm{p}_{\text{target}}]$. The optimization objective is to find $(\bm{q}^{*},\bm{d}^{*})$ minimizing the weighted position/orientation errors (with consistent units, e.g., meters for position, radians for orientation):
@@
where $\bm{p}\in\mathbb{R}^{3}$ and $\bm{R}\in SO(3)$ are the actual end-effector position and orientation, $\bm{p}_{\text{target}}$ and $\bm{R}_{\text{target}}$ are the desired pose components, $d_{R}$ is the Riemannian distance on $SO(3)$, defined as $d_{R}(\bm{R}_{1},\bm{R}_{2})=\arccos((\mathop{\mathrm{Tr}}\nolimits(\bm{R}_{1}^{\top}\bm{R}_{2})-1)/2)$ in radians, $\bm{d}\in\mathcal{D}$ denotes design parameters, and $\alpha_{p},\alpha_{R}\geq 0$ are position and orientation error weights, respectively, chosen to balance units (e.g., $\alpha_{p}=1$ for meters squared, $\alpha_{R}=1$ for radians squared, or scaled empirically based on task requirements). To address the combinatorial complexity inherent in high-dimensional optimization, we reformulate the problem into a quantum-native representation using the following computational framework, considering NISQ hardware limitations such as noise and limited qubits.
@@ III-B Quantum Computing Framework
The transformation from classical kinematics to quantum computation is achieved through five integrated stages. To make this accessible to robotics experts, we provide explanations of quantum concepts, analogies to robotic systems (e.g., quantum superposition akin to parallel evaluation of multiple manipulator configurations), and references to key literature.
III-B1 Step 1: Discretization & Quantum State Encoding
@@
For angular parameters like joint angles, the bounds $z_{i}^{\min},z_{i}^{\max}]=[0,2\pi)$ account for periodicity, with wrap-around handled in decoding to avoid discontinuities. This discretization is similar to gridding the configuration space in robotic path planning, where continuous joint angles are sampled at finite resolutions to make the search tractable. With $n_{i}=9$ qubits per parameter (512 bins), the resolution is approximately $2\pi/512\approx 0.012$ radians ($\approx 0.7^{\circ}$) for angles and similar for linear parameters (e.g., link lengths in [0.1, 2.0] m yield $\approx 0.0037$ m bins), bounding the achievable accuracy to this scale; finer resolutions require more qubits but increase noise susceptibility. The full configuration space is encoded into $N=\sum n_{i}$ qubits via uniform superposition:
@@
where $\lvert k\rangle$ represents the quantum state corresponding to the discretized parameter vector $\bm{z}^{(k)}$. [^1] The bra-ket (Dirac) notation denotes quantum states: $\lvert\psi\rangle$ (ket) is a column vector in a complex Hilbert space, analogous to a state vector in robotics kinematics and dynamics; $\langle\psi\rvert$ (bra) is its conjugate transpose. Inner products $\langle\psi\rvert\lvert\phi\rangle$ measure similarity, and $\lvert k\rangle$ here encodes a specific robotic configuration as a basis state. A qubit can be regarded as an abstract robotic parameter in superposition, enabling simultaneous representation of multiple kinematic configurations. The uniform superposition $\lvert\psi\rangle$ is analogous to a robotic swarm exploring distinct configurations in parallel, offering simultaneous evaluation without sequential computation—a key quantum advantage.
III-B2 Step 2: Quantum Machine Learning for Kinematic Approximation

### v2

@@ Abstract
Optimizing high-degree-of-freedom robotic manipulators requires searching complex, high-dimensional configuration spaces, a task that is computationally challenging for classical methods. This paper introduces a quantum-native framework that integrates Quantum Machine Learning (QML) with Grover’s algorithm to solve kinematic optimization problems efficiently. A parameterized quantum circuit is trained to approximate the forward kinematics model, which then constructs an oracle to identify optimal configurations. Grover’s algorithm leverages this oracle to provide a quadratic reduction in search complexity. Demonstrated on simulated 1-DoF, 2-DoF, and dual-arm manipulator tasks, the method achieves significant speedups—up to 93x over classical optimizers like Nelder-Mead—as problem dimensionality increases. This work establishes a foundational, quantum-native framework for robot kinematic optimization, effectively bridging quantum computing and robotics problems.
@@ I Introduction
Robotic manipulators are pivotal in modern automation, manufacturing, and precision tasks across industries including automotive assembly, medical surgery, and aerospace [1, 2, 3]. These systems require precise end-effector positioning, which is achieved through inverse kinematics (IK) to determine the joint configurations corresponding to the target poses [4, 5]. Classical IK optimization methods perform adequately for low-DoF manipulators with simple workspaces but encounter limitations in high-dimensional, nonlinear manifolds [6, 7]. In parallel robot optimization or design candidate selection, challenges intensify due to nonlinear constraints, high-dimensional spaces, and mixed discrete-continuous parameters [8]. Exhaustive searches over discretized joints scale exponentially as $O(k^{N})$ for $N$-DoF systems with $k$
@@
values per joint, while advanced algorithms exhibit polynomial scaling, becoming intractable for high-DoF manipulators.
This computational bottleneck motivates quantum computing, leveraging superposition and entanglement for parallel processing and asymptotic advantages in search and optimization [9, 10, 11]. Grover’s algorithm searches unsorted databases of $M$ elements in $O(\sqrt{M})$ queries, offering quadratic speedup over classical $O(M)$ methods [12]. In high-DoF robot IK optimization where $M\sim k^{N}$, this enables efficient exploration of constrained configuration spaces. Quantum approaches handle exponential complexity in non-convex landscapes, advancing robotics beyond classical limits as hardware matures. Consequently, quantum computing applications in robotics are expanding across multiple fronts, including Grover’s algorithm for optimal control [13], quantum annealing for IK via QUBO formulation with 30x speedups [14], QML for robot test oracles [15], quantum-inspired reinforcement learning for control [16], Grover-based classification [17], quantum-inspired sliding-mode control [18], and quantum deep RL for navigation [19].
However, existing approaches exhibit fragmented quantum-classical pipelines. For instance, while [13] uses Grover’s search standalone and [14] relies on specialized annealing hardware, none integrate QML for learning with Grover’s search within a unified quantum-native framework—an end-to-end pipeline embedded entirely in quantum circuits.
To address this, we propose a novel quantum framework integrating QML for learning kinematic patterns using parameterized quantum circuits with Grover’s search for efficient optimization. This quantum-native approach avoids classical-quantum data transfer overhead during computation. As shown in Fig. 1, our framework comprises a QML block for state preparation and a Grover block where an oracle marks low-cost states based on observables (Table I).
@@
Note: Observables are diagonal in computational basis for quantum-parallel evaluation. $\bm{p}_{k}$, $\bm{p}_{\text{target}}$: end-effector and target positions; $\bm{R}_{k}$, $\bm{R}_{\text{target}}$: orientation matrices, $d_{R}$: SO(3) geodesic; $\bm{J}_{k}$: Jacobian; $\bm{\tau}_{k}$: torques; $d_{\text{min}}(\bm{q}_{k})$: minimum obstacle distance. Projector $\lvert k\rangle\langle k\rvert$ maps configuration $k$ to quantum state.
We demonstrate the feasibility of the proposed approach through implemented basic examples. With the continued advancement of scalable, error-corrected quantum processors [20], this method shows strong potential to tackle higher-dimensional robotic problems that are exceedingly difficult to solve using classical computation.
@@
The paper is organized as follows: Section II presents the basics of quantum computing for robotics, describes its applications in robotics, and reviews related work; Section III details the framework; Section IV describes case studies; Section V presents evaluations; Section VI concludes.
II Background and Related Works
@@
This section surveys quantum computing advances in robotics, explaining key concepts for accessibility: qubits, superposition, entanglement, and bra-ket notation.
The motion of robotic manipulators is fundamentally governed by kinematics [3]. Forward kinematics (FK) maps the joint variables $\bm{q}\in\mathbb{R}^{N}$ to the end-effector pose $\bm{x}\in SE(3)$, whereas IK performs the inverse mapping $\bm{q}=f^{-1}(\bm{x})$. These problems are inherently nonlinear and multi-valued for high-DoF or redundant systems ($N>6$), often formulated as a high-dimensional optimization task.
Classical geometric and metaheuristic solvers suffer from scalability limitations arising from local minima, singularities (i.e., rank deficiency of the Jacobian $\bm{J}(\bm{q})$), and most importantly exponential computational growth in high-DoF or constrained environments [5]. These challenges motivate the exploration of quantum paradigms that leverage intrinsic parallelism to achieve global optimization.
II-A Quantum Computing Essentials for Robotics
Quantum computers operate on qubits with states $\lvert\psi\rangle=\alpha\lvert 0\rangle+\beta\lvert 1\rangle$ ($\alpha,\beta\in\mathbb{C}$, $|\alpha|^{2}+|\beta|^{2}=1$) enabling superposition of multiple robotic configurations. Entanglement correlates qubits for joint operations. Bra-ket notation: $\lvert\cdot\rangle$ denotes column vectors, $\langle\cdot\rvert$ conjugate transposes; projectors like $\lvert k\rangle\langle k\rvert$ isolate states. Quantum circuits apply unitary gates (Hadamard for superposition, CNOT for entanglement). Grover’s algorithm provides a quadratic speedup through amplitude amplification, which operates in two steps per iteration. First, the oracle $O_{f}$ identifies solution states by applying a phase inversion ($\lvert k\rangle\mapsto-\lvert k\rangle$); second, the diffusion operator $D=2\lvert\psi_{0}\rangle\langle\psi_{0}\rvert-I$ reflects amplitudes about their mean, amplifying the probability amplitudes of feasible manipulator configurations. After approximately $\mathcal{O}(\sqrt{N/M})\approx\frac{\pi}{4}\sqrt{N/M}$ iterations, the probability of measuring an optimal configuration becomes maximal.
II-B Quantum Applications in Robotics and Related Domains
@@
Quantum methods have been applied to various robotics challenges. Some of the recent works include: Dahassa et al. [13, 21], who reformulated optimal control as Grover’s search and quantum genetic algorithms for six-jointed arms, utilizing quantum comparators and sigmoid updates for convergence in classical simulation; Salloum et al. [14], who expressed IK as a Quadratic Unconstrained Binary Optimization (QUBO) problem for quantum annealing, achieving up to 30× speedups through hybrid solvers; Wang et al. [15], who employed quantum reservoir computing for robot test oracles, reducing errors by 15%; Yu et al. [16] and Kirchner et al. [19], who utilized quantum-inspired reinforcement learning for control and navigation; and Du et al. [17], who enhanced Grover-based classification with fewer measurements. Other emerging applications include quantum teleportation for remote robot control [22], quantum-enhanced social robotics [23], quantum-inspired sliding-mode control for industrial arms [18], and hybrid LLM-Q-learning for adaptive robotics [24].
Unlike these fragmented approaches, our framework integrates QML for learning kinematic representations with Grover’s algorithm for quantum search, forming a fully quantum-native optimization pipeline. A more complete mapping of application domains to quantum and quantum-inspired algorithms is provided in Table II.
@@
II-C Overview of the Proposed Framework
Our framework encodes kinematic parameters $\bm{z}\in\mathbb{R}^{d}$ into parameterized circuit $U_{\text{QML}}(\bm{\theta})$ approximating forward kinematics $f(\bm{z})$. Data encoding uses rotation gates ($R_{X},R_{Y},R_{Z}$); training minimizes cost $L(\bm{z})=\|f(\bm{z})-\bm{p}_{\text{target}}\|^{2}$. For Grover’s, discretize $\bm{z}$ with $n_{i}$ qubits per parameter, yielding $N=\sum n_{i}$ qubits and space $\mathbb{C}^{2^{N}}$. Basis states $\lvert x\rangle$ map to configurations $\bm{z}_{x}$. The oracle marks low-cost states; Grover amplifies them for near-optimal solutions, exploiting QML’s representation and Grover’s efficiency in a quantum-native approach.
@@ III-A Problem Statement: Robotic Kinematics Optimization
When geometric and joint parameters are unknown but the task is specified, optimization identifies parameters $\bm{d}$ and joint configuration $\bm{q}$ enabling task achievement. For each candidate $\bm{d}$, inverse kinematics checks existence of $\bm{q}$ achieving $\bm{T}_{\text{target}}$. If infeasible or error exceeds tolerance, $\bm{d}$ is updated iteratively. This combines IK with parameter adjustment, ensuring the final design achieves $\bm{T}_{\text{target}}$ within kinematic limits. Formally: Consider an $n$-DoF manipulator with $\bm{q}=[q_{1},\dots,q_{n}]^{\top}$, forward kinematics $f_{\text{FK}}:\mathbb{R}^{n}\times\mathcal{D}\rightarrow SE(3)$ parameterized by $\bm{d}\in\mathcal{D}$, and target pose $\bm{T}_{\text{target}}\in SE(3)$ represented as a 4×4 homogeneous transformation matrix $\bm{T}_{\text{target}}=\begin{bmatrix}\bm{R}_{\text{target}}&\bm{p}_{\text{target}}\\ \bm{0}^{\top}&1\end{bmatrix}$.
The optimization objective is to find $(\bm{q}^{*},\bm{d}^{*})$ minimizing the weighted position/orientation errors:
@@
where $\bm{p}\in\mathbb{R}^{3}$ and $\bm{R}\in SO(3)$ are the actual end-effector position and orientation, $\bm{p}_{\text{target}}$ and $\bm{R}_{\text{target}}$ are the desired pose components, $d_{R}$ is the Riemannian distance on $SO(3)$, defined as $d_{R}(\bm{R}_{1},\bm{R}_{2})=\arccos((\mathop{\mathrm{Tr}}\nolimits(\bm{R}_{1}^{\top}\bm{R}_{2})-1)/2)$ in radians, $\bm{d}\in\mathcal{D}$ denotes design parameters, and $\alpha_{p},\alpha_{R}\geq 0$ are position and orientation error weights, respectively, chosen to balance units (e.g., $\alpha_{p}=1$ for meters squared, $\alpha_{R}=1$ for radians squared, or scaled empirically based on task requirements). To address the combinatorial complexity inherent in high-dimensional optimization, we reformulate the problem into a quantum-native representation using the following computational framework, considering Noisy Intermediate-Scale Quantum (NISQ) hardware limitations such as noise and limited qubits.
@@ III-B Quantum Computing Framework
The transformation from classical kinematics to quantum computation is achieved through the following five integrated stages.
III-B1 Discretization & Quantum State Encoding
@@
For angular parameters like joint angles, the bounds $z_{i}^{\min},z_{i}^{\max}]=[0,2\pi]$ account for periodicity, with wrap-around handled in decoding to avoid discontinuities. This discretization is similar to gridding the configuration space in robotic path planning, where continuous joint angles are sampled at finite resolutions to make the search tractable. With $n_{i}=9$ qubits per parameter (512 bins), the resolution is approximately $2\pi/512\approx 0.012$ radians ($\approx 0.7^{\circ}$) for angles and similar for linear parameters (e.g., link lengths in [0.1, 2.0] m yield $\approx 0.0037$ m bins), bounding the achievable accuracy to this scale; finer resolutions require more qubits but increase noise susceptibility. This choice balances resolution with NISQ hardware constraints (limited qubits and coherence times). The full configuration space is encoded into $N=\sum n_{i}$ qubits via uniform superposition:
@@
where $\lvert k\rangle$ represents the quantum state corresponding to the discretized parameter vector $\bm{z}^{(k)}$. [^1] The bra-ket (Dirac) notation denotes quantum states: $\lvert\psi\rangle$ (ket) is a column vector in a complex Hilbert space, analogous to a state vector in robotics kinematics and dynamics; $\langle\psi\rvert$ (bra) is its conjugate transpose. Inner products $\langle\psi\rvert\lvert\phi\rangle$ measure similarity, and $\lvert k\rangle$ here encodes a specific robotic configuration as a basis state. A qubit can be regarded as an abstract robotic parameter in superposition, enabling simultaneous representation of multiple kinematic configurations. The uniform superposition $\lvert\psi\rangle$ is analogous to a robotic swarm exploring distinct configurations in parallel, offering simultaneous evaluation without sequential computation.
III-B2 QML for Kinematic Approximation